# From Confidence to Collapse in LLM Factual Robustness

### v2

@@ Abstract
Ensuring the robustness of factual knowledge in LLMs is critical for reliable applications in tasks such as question answering and reasoning. However, existing evaluation methods predominantly focus on performance-based metrics, often investigating from the perspective of prompt perturbations, which captures only the externally triggered side of knowledge robustness. To bridge this gap, we introduce a principled approach to measure factual robustness from the perspective of the generation process by analyzing token distribution entropy in combination with temperature scaling sensitivity. These two factors build the Factual Robustness Score (FRS), a novel metric which quantifies the stability of a fact against perturbations in decoding conditions, given its initial uncertainty. To validate our approach, we conduct extensive experiments on 5 LLMs across 3 closed-book QA datasets (SQuAD, TriviaQA, and HotpotQA). We show that factual robustness varies significantly – smaller models report an FRS of $0.76$, larger ones $0.93$ – with accuracy degrading by $\sim$$60\%$ under increased uncertainty. These insights demonstrate how entropy and temperature scaling impact factual accuracy, and lay a foundation for developing more robust knowledge retention and retrieval in future models.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized natural language understanding and generation, demonstrating remarkable capabilities across tasks such as question answering, knowledge-intensive text generation, and reasoning (Petroni et al., 2019; Roberts et al., 2020; Hu et al., 2023). However, despite their strong performance, factual stability remains an open challenge. While a model may provide a correct answer under one set of decoding conditions, it may fail to do so when faced with minor perturbations, such as variations in sampling temperature. This raises a fundamental question: how robustly is factual knowledge embedded in LLMs?
Traditional evaluations of factual knowledge in LLMs rely on accuracy-based metrics, which measure correctness under fixed conditions (Joshi et al., 2017; Kwiatkowski et al., 2019; Hendrycks et al., 2021). However, these methods do not account for uncertainty in fact retrieval – whether a model is inherently confident in its answer or if its correctness is fragile under perturbations. Additionally, prior techniques often focus on single-temperature evaluations, failing to systematically analyze how factual outputs degrade as uncertainty is introduced (Maynez et al., 2020; Gabriel et al., 2021). To bridge this gap, we introduce a novel perspective: factual robustness, which examines not just whether an answer is correct but how resistant it is to internal uncertainty and temperature-induced variability.
To quantify this, we propose the Factual Robustness Score (FRS), a new metric designed to measure the stability of factual knowledge within an LLM. Unlike previous work that assesses model performance at isolated temperature values, FRS integrates both entropy and temperature, providing a comprehensive robustness assessment. Entropy quantifies the model’s intrinsic confidence in an answer, while the so-called “breaking” temperature captures how much uncertainty a fact can withstand before becoming incorrect. By combining these two dimensions, FRS moves beyond simple accuracy and provides a deeper insight into how reliably knowledge is stored and retrieved within LLMs.
@@ 2 Preliminaries
Temperature is a parameter that controls the sharpness of the probability distribution during text generation. It operates by scaling the logits before applying the softmax function, as defined in Equation˜2:
@@ 3.1 Factual Knowledge in Language Models
LLMs encode vast amounts of factual knowledge, often retrieved from internal representations rather than structured databases. While models can recall these facts, their responses remain inconsistent, suffering from hallucinations and retrieval errors (Petroni et al. (2019), Roberts et al. (2020)). A prominent research direction focuses on understanding and modifying how knowledge is encoded within LLMs: while Geva et al. (2020) and Dai et al. (2021) identified specific structures within Transformers, such as key-value memory stores and knowledge neurons, De Cao et al. (2021) and Meng et al. (2022) explored techniques to directly manipulate the information stored within model weights. However, beyond the challenge of modifying knowledge, ensuring its stability and reliability remains an open problem. While Zong et al. (2024) introduced ComparisonQA to evaluate factual robustness under controlled knowledge frequency and uncertainty, a broader body of work has framed robustness through the lens of adversarial attacks. Studies such as Xu et al. (2024), Bondarenko and Viehweger (2024), and Howe et al. (2024) have examined the vulnerability of LLMs to prompt-side perturbations, revealing inconsistencies in factual recall. Approaching the question through the prompt lens, Mahaut et al. (2024) consider factual robustness in terms of semantically equivalent question phrasing. Contrary to the approaches above, our work shifts the focus from external prompt modifications to the model-side of the generation process, investigating the internal mechanisms that contribute to factual robustness.
@@ 3.2 Temperature and Language Models
While seemingly distant from modern LLMs, Ackley et al. (1985) established temperature as a key factor in shaping probability distributions, a principle still central to language model decoding today. Recent studies have debated whether temperature significantly impacts problem-solving ability. Renze and Guven (2024) found that varying temperature from 0.0 to 1.6 had no significant effect on accuracy, challenging the assumption that lower temperatures enhance reasoning. Beyond accuracy, temperature is often associated with creativity: Peeperkorn et al. (2024) showed that while higher temperatures increase variation, this does not necessarily translate into meaningful novelty, raising questions about its role.
Recognizing the limitations of fixed-temperature sampling, researchers have developed adaptive temperature control strategies to balance quality and diversity. Chang et al. (2023) introduced KL-Divergence Guided Temperature Sampling, which adjusts temperature dynamically based on token relevance, while Zhang et al. (2024) proposed Entropy-based Dynamic Temperature (EDT) Sampling to optimize fluency and diversity.
A novel perspective on temperature scaling comes from Nakaishi et al. (2024), who found that LLMs undergo critical phase transitions at certain temperature thresholds, exhibiting abrupt behavioral shifts akin to phase changes in physical systems. This suggests that temperature adjustments do more than refine probability distributions – they can fundamentally reshape generative dynamics, making temperature a key factor in understanding LLM computational behavior.
@@
Unlike existing performance-based methods that assess correctness at fixed temperatures, or studies in attack-based scenarios, we evaluate how robustly a fact is stored within a model by incorporating both entropy and “breaking” temperature, a temperature where the models lose answer confidence. This allows for a more comprehensive assessment of factual stability, moving beyond surface-level correctness to quantify how knowledge withstands perturbations in sampling conditions. Our work provides new insights into the resilience of LLMs and highlights the need for robustness-focused evaluation metrics in future research.
@@ 4.1.1 Closed-Book Question Answering Setup
Datasets. We use three widely studied QA datasets, each adjusted for the closed-book setting: (1) SQuAD (Rajpurkar et al., 2016), a reading comprehension dataset primarily composed of Wikipedia-based questions; (2) TriviaQA (Joshi et al., 2017), a collection of knowledge-intensive questions originally designed for open-domain QA; and (3) HotpotQA (Yang et al., 2018), a multi-hop QA dataset that requires reasoning over multiple facts.
Model Selection. To ensure diversity in model architectures, scales, and training sources, we evaluate five different LLMs: i.e., GPT-4o-mini (OpenAI, 2023), LLaMA-3.2-3B and LLaMA-3.1-8B (Touvron et al., 2023), Qwen-2.5-3B and Qwen-2.5-14B (Yang et al., 2024). The GPT models are accessed via the OpenAI API, while the LLaMA and Qwen models are loaded from Hugging Face, using their instruct fine-tuned versions.
@@ 4.1.2 Selection of Correctly Answered Questions for Analysis
Baseline Selection Criteria. To identify these correctly answered questions, we set the generation temperature $t$ to zero, enforcing deterministic output selection. This choice minimizes stochastic variability, ensuring that the most probable token sequence is always selected. By setting $t=0$, we establish a baseline representing the model’s most “ideal” conditions for factual retrieval. Additionally, to simulate a well-calibrated model, we filter out instances where the model confidently produces incorrect answers. This ensures that our study focuses solely on high-confidence correct responses, allowing us to analyze the impact of increasing temperature without confounding effects from miscalibrated predictions.
@@ 4.2 Temperature Scaling
Temperature $t$ is a key parameter in the generation process, controlling the level of randomness and semantic variability in model outputs. It theoretically ranges from 0 to $\infty$, although it is typically constrained between 0 and $2$ (Renze and Guven, 2024). Recall that lower temperatures lead to more deterministic outputs, while higher values introduce greater diversity by flattening the probability distribution. To systematically analyze the impact of temperature on response accuracy and certainty, we apply temperature scaling to the previously identified correctly answered questions for each model and dataset. The procedure is as follows:
@@ 4.3 Entropy of true distribution
Besides temperature, we analyze the entropy levels of generated answers to understand how entropy interacts with temperature-induced distribution shifts without compromising factual accuracy. To standardize our entropy measurements, we use $\log_{10}$ instead of $\log_{2}$, ensuring that the maximum entropy value is fixed at 1. The theoretical upper bound of entropy for a given probability distribution is determined by $\log(n)$, where $n$ represents the number of possible outcomes – in our case, the number of candidate tokens at a given position. Since we restrict the top token choices to 10, the maximum entropy value in our setup is $\log_{10}(10)=1$.
@@ 5 How Factually Robust are LLMs?
In contrast to the findings of Renze and Guven (2024), who argue that temperature has no significant impact, our results in Sections˜4.2 and 4.3 demonstrate that increasing temperature in a QA context reduces both accuracy and certainty, regardless of model size. This sensitivity to temperature highlights the need for LLMs to maintain factual reliability under varying generation conditions.
@@ 5.1 Factual Robustness Score (FRS)
Figure 4: FRS ($d=1$) over all models, on SQuAD. Although FRS equal to 1 is theoretically achievable when $t_{b}\to\infty$, in practice, we set it to 1 for all samples where models did not break with $t_{b}\leq 2$. Hence, the yellow boxes.
@@
Table 1: Average FRS with $d\in\{1,2,5,10,50\}$ across all datasets. Generally larger models are more robust than smaller ones; however the difference is not significant. indicates the most robust model overall.
@@
Table˜1 compares FRS across different model sizes and families using various values of $d\in\{1,2,5,10,50\}$. As expected, larger models tend to be more robust than their smaller counterparts, although not significantly. A particularly interesting finding is that model size alone is not the best predictor of robustness across different model families. For example, even though LLaMA-8b exhibits a more stable entropy trend in relation to $t_{b}$ (see Figure˜3), the smaller Qwen-3b model actually achieves a higher average FRS. We argue that different models store and retrieve factual knowledge in distinct ways. While larger models within the same family (e.g., LLaMA-3b vs. LLaMA-8b) tend to retain facts more reliably, comparing across families (e.g., LLaMA vs. Qwen) introduces additional factors beyond just model size. We also show how the FRS gets hampered with an increasing $d$ based on the initial entropy. Notice how the LLaMA models suffer the highest drop in FRS (0.174-0.174 for 3b, and 0.149-0.149 for 8b), meaning that they were uncertain, on average, about the initial factual knowledge, confirmed in Figure˜1 where LLaMA has the highest accuracy drop across the board. Most surprisingly, GPT-4o-mini is either less robust or, at best, on par with Qwen-14b when $d\to\infty$. This adds another dimension to the ongoing debate about whether smaller, more efficient LLMs that require fewer computing resources can outperform OpenAI’s large, closed-source models. To explore this further, Figure˜4 illustrates FRS with $d=1$ across all models on the SQuAD dataset. Notably, Qwen-14b achieves higher FRS scores even when its initial entropy is lower than that of GPT-4o-mini, whereas the latter’s scores are more concentrated toward the lower entropy range. Additionally, GPT-4o-mini exhibits a more skewed probability density distribution, with $\text{FRS}\rightarrow 1$, suggesting that while it may be highly confident in certain cases, its robustness is not consistently superior across all factual scenarios (see Appendix˜G). These findings highlight the need for a deeper examination into how architectures, training methodologies, and knowledge retention influence factual robustness.
@@ 5.3 Most and Least Robust Facts
We further analyze the specific question types that are the most and least robust within the models. For this, we choose to fine-tune a lightweight classifier, DistilBERT (Sanh et al., 2019), on the TREC question type dataset (Li and Roth, 2002) to classify the categories Numerical, Location, Entity and Human, for which we reach 96,6% accuracy on the test set. [^3] We report details about the dataset and the training procedure in Appendix H. We then collect the top 1000 highest and lowest FRS score samples across all datasets, for each model. For both the top and bottom samples, we compute the proportion of each question type relative to the total number of samples of that type. Figure˜5 shows the ratio of the most robust question types in percentage for each model. For example, for LLaMA-3B, 64.8% of numerical type questions are among the most robust (while $100\%-64.8\%=35.2\%$ are part of the least robust). Answers associated with numerical and location related questions tend to be the most robust across models, reaching up to 72.9% and 64.9%, respectively. The least robust answers belong to the Human entity type, which means questions asking about the name of a person or a group. We show that all models recall names in a less robust way, meaning that those are most prone to be produced with high entropy and/or are shifting to a wrong answer at low temperature levels. We hypothesize that this is due to the nature of expected answers: for example, questions of the numerical type most often call for a single digit as answer, e.g., a year. Here, the LLM has less individual tokens to produce, and hence less chance of introducing an error, than for example with the name of a person, which is more lengthy. It is furthermore interesting to note the difference between model sizes: while the bigger models perform highest on numerical type questions, they underperform smaller models like LLaMA-3B on human type questions.

### v3

@@ Abstract
Ensuring the robustness of factual knowledge in LLMs is critical for reliable applications in tasks such as question answering and reasoning. However, existing evaluation methods predominantly focus on performance-based metrics, often investigating from the perspective of prompt perturbations, which captures only the externally triggered side of knowledge robustness. To bridge this gap, we introduce a principled approach to measure factual robustness from the perspective of the generation process by analyzing token distribution entropy in combination with temperature scaling sensitivity. These two factors build the Factual Robustness Score (FRS), a novel metric which quantifies the stability of a fact against perturbations in decoding conditions, given its initial uncertainty. To validate our approach, we conduct extensive experiments on 5 LLMs across 3 closed-book QA datasets (SQuAD, TriviaQA, and HotpotQA). We show that factual robustness varies significantly – smaller models report an FRS of $0.76$, larger ones $0.93$ – with accuracy degrading by $\sim$$60\%$ under increased uncertainty. These insights demonstrate how entropy and temperature scaling impact factual accuracy, and lay a foundation for developing more robust knowledge retention and retrieval in future models. We release our code at https://github.com/afastowski/frs.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized natural language understanding and generation, demonstrating remarkable capabilities across tasks such as question answering, knowledge-intensive text generation, and reasoning (Petroni2019Knowledge; Roberts2020HowMuch; hu2023large). However, despite their strong performance, factual stability remains an open challenge. While a model may provide a correct answer under one set of decoding conditions, it may fail to do so when faced with minor perturbations, such as variations in sampling temperature. This raises a fundamental question: how robustly is factual knowledge embedded in LLMs?
Traditional evaluations of factual knowledge in LLMs rely on accuracy-based metrics, which measure correctness under fixed conditions (2017arXivtriviaqa; kwiatkowski2019natural; hendrycks2020measuring). However, these methods do not account for uncertainty in fact retrieval – whether a model is inherently confident in its answer or if its correctness is fragile under perturbations. Additionally, prior techniques often focus on single-temperature evaluations, failing to systematically analyze how factual outputs degrade as uncertainty is introduced (Maynez2020Faithfulness; gabriel2021go). To bridge this gap, we introduce a novel perspective: factual robustness, which examines not just whether an answer is correct but how resistant it is to internal uncertainty and temperature-induced variability.
To quantify this, we propose the Factual Robustness Score (FRS), a new metric designed to measure the stability of factual knowledge within an LLM. Unlike previous work that assesses model performance at isolated temperature values, FRS integrates both entropy and temperature, providing a comprehensive robustness assessment. Entropy quantifies the model’s intrinsic confidence in an answer, while the so-called “breaking” temperature is the first temperature greater than zero, at which the generation produces an incorrect answer. In other words, it captures how much uncertainty the generation process can withstand, before shifting from the correctly produced answer to an incorrect one. By combining these two dimensions, FRS moves beyond simple accuracy and provides a deeper insight into how reliably knowledge is stored and retrieved within LLMs.
@@ 2 Preliminaries
Temperature is a parameter that controls the sharpness of the probability distribution during text generation. It operates by scaling the logits before applying the softmax, as defined in Equation˜2:
@@ 3.1 Factual Knowledge in Language Models
LLMs encode vast amounts of factual knowledge, often retrieved from internal representations rather than structured databases. While models can recall these facts, their responses remain inconsistent, suffering from hallucinations and retrieval errors (Petroni2019Knowledge, Roberts2020HowMuch). A prominent research direction focuses on understanding and modifying how knowledge is encoded within LLMs: while geva2020transformer and dai2021knowledge identified specific structures within Transformers, such as key-value memory stores and knowledge neurons, de2021editing and meng2022locating explored techniques to directly manipulate the information stored within model weights. However, beyond the challenge of modifying knowledge, ensuring its stability and reliability remains an open problem. While zong2024comparisonqa introduced ComparisonQA to evaluate factual robustness under controlled knowledge frequency and uncertainty, a broader body of work has framed robustness through the lens of adversarial attacks. Studies such as xu2024llmfoolitself, bondarenko2024llm, and howe2024exploring have examined the vulnerability of LLMs to prompt-side perturbations, revealing inconsistencies in factual recall. Approaching the question through the prompt lens, mahaut2024factual consider factual robustness in terms of semantically equivalent question phrasing. Contrary to the approaches above, our work shifts the focus from external prompt modifications to the model-side of the generation process, investigating the internal mechanisms that contribute to factual robustness.
@@ 3.2 Temperature and Language Models
While seemingly distant from modern LLMs, ackley1985learning established temperature as a key factor in shaping probability distributions, a principle still central to language model decoding today. Recent studies have debated whether temperature significantly impacts problem-solving ability. Renze24Sampling found that varying temperature from 0.0 to 1.6 had no significant effect on accuracy, challenging the assumption that lower temperatures enhance reasoning. Beyond accuracy, temperature is often associated with creativity: peeperkorn2024temperature showed that while higher temperatures increase variation, this does not necessarily translate into meaningful novelty, raising questions about its role.
Recognizing the limitations of fixed-temperature sampling, researchers have developed adaptive temperature control strategies to balance quality and diversity. chang2023kl introduced KL-Divergence Guided Temperature Sampling, which adjusts temperature dynamically based on token relevance, while zhang2024edt proposed Entropy-based Dynamic Temperature (EDT) Sampling to optimize fluency and diversity.
A novel perspective on temperature scaling comes from nakaishi2024critical, who found that LLMs undergo critical phase transitions at certain temperature thresholds, exhibiting abrupt behavioral shifts akin to phase changes in physical systems. This suggests that temperature adjustments do more than refine probability distributions – they can fundamentally reshape generative dynamics, making temperature a key factor in understanding LLM computational behavior.
@@
Unlike existing performance-based methods that assess correctness at fixed temperatures, or studies in attack-based scenarios, we evaluate how robustly a fact is stored within a model by incorporating both entropy and “breaking” temperature – the temperature at which a model switches to producing an incorrect answer. This allows for a more comprehensive assessment of factual stability, moving beyond surface-level correctness to quantify how knowledge withstands perturbations in sampling conditions. Our work provides new insights into the resilience of LLMs and highlights the need for robustness-focused evaluation metrics in future research.
@@ 4.1.1 Closed-Book Question Answering Setup
Datasets. We use three widely studied QA datasets, each adjusted for the closed-book setting: (1) SQuAD (rajpurkar-etal-2016-squad), a reading comprehension dataset primarily composed of Wikipedia-based questions; (2) TriviaQA (2017arXivtriviaqa), a collection of knowledge-intensive questions originally designed for open-domain QA; and (3) HotpotQA (yang2018hotpotqa), a multi-hop QA dataset that requires reasoning over multiple facts.
Model Selection. To ensure diversity in model architectures, scales, and training sources, we evaluate five different LLMs: i.e., GPT-4o-mini (achiam2023gpt), LLaMA-3.2-3B and LLaMA-3.1-8B (touvron2023llama), Qwen-2.5-3B and Qwen-2.5-14B (yang2024qwen2). The GPT models are accessed via the OpenAI API, while the LLaMA and Qwen models are loaded from Hugging Face, using their instruct fine-tuned versions.
@@ 4.1.2 Selection of Correctly Answered Questions for Analysis
Baseline Selection Criteria. To identify these correctly answered questions, we set the generation temperature $t$ to zero, enforcing deterministic output selection. This choice minimizes stochastic variability, ensuring that the most probable token sequence is always selected. Since we study model robustness, we are showing the effects of increasingly hard generation conditions. Hence, by first setting $t=0$, we establish a baseline representing the model’s most “ideal” conditions for factual retrieval. Additionally, to simulate a well-calibrated model, we filter out instances where the model confidently produces incorrect answers. This ensures that our study focuses solely on high-confidence correct responses, allowing us to analyze the impact of increasing temperature without confounding effects from miscalibrated predictions.
@@ 4.2 Temperature Scaling
Temperature $t$ is a key parameter in the generation process, controlling the level of randomness and semantic variability in model outputs. It theoretically ranges from 0 to $\infty$, although it is typically constrained between 0 and $2$ (Renze24Sampling). Recall that lower temperatures lead to more deterministic outputs, while higher values introduce greater diversity by flattening the probability distribution. To systematically analyze the impact of temperature on response accuracy and certainty, we apply temperature scaling to the previously identified correctly answered questions for each model and dataset. The procedure is as follows:
@@ 4.3 Entropy of true distribution
Besides temperature, we analyze the entropy levels of generated answers to understand how entropy interacts with temperature-induced distribution shifts without compromising factual accuracy. To standardize our entropy measurements, we use $\log_{10}$ instead of $\log_{2}$, ensuring that the maximum entropy value is fixed at 1. The theoretical upper bound of entropy for a given probability distribution is determined by $\log(n)$, where $n$ represents the number of possible outcomes – in our case, the number of candidate tokens at a given position. Since we restrict the top token choices to 10222The restriction is necessary due to API limitations for GPT-4o-mini. In order to be consistent, we apply the same limit to all models., the maximum entropy value in our setup is $\log_{10}(10)=1$.
@@ 5 How Factually Robust are LLMs?
In contrast to the findings of Renze24Sampling, who argue that temperature has no significant impact, our results in Sections˜4.2 and 4.3 demonstrate that increasing temperature in a QA context reduces both accuracy and certainty, regardless of model size. This sensitivity to temperature highlights the need for LLMs to maintain factual reliability under varying generation conditions.
@@ 5.1 Factual Robustness Score (FRS)
Figure 4: FRS ($d=1$) over all models, on SQuAD. Although FRS equal to 1 is theoretically achievable when $t_{b}\to\infty$, in practice, we set it to 1 for all samples where models did not break with $t_{b}\leq 2$. Hence, the yellow data points.
@@
Table 1: Average FRS with $d\in\{1,2,5,10,50\}$ across all datasets. Generally larger models are more robust than smaller ones. indicates the most robust model overall.
@@
Table˜1 compares FRS across different model sizes and families using various values of $d\in\{1,2,5,10,50\}$, controlling the influence of the entropy. As expected, larger models generally tend to be more robust than their smaller counterparts. However, a particularly interesting finding is that model size alone is not the sole predictor of robustness across different model families. While larger models within the same family (e.g., LLaMA-3b vs. LLaMA-8b) tend to retain facts more reliably, comparing across families (e.g., LLaMA vs. Qwen) introduces additional factors beyond just model size. Even though LLaMA-8b exhibits a more stable entropy trend in relation to $t_{b}$ (see Figure˜3), the smaller Qwen-3b model actually achieves a higher average FRS. Interesting in this context is also GPT-4o-mini, which, being an 8 billion parameters model, has similarly high robustness as the leader, Qwen-14b. Hence, we argue that different models store and retrieve factual knowledge in distinct ways. While parameter size plays a central role in a model’s factual robustness, factors like model architecture and training data may also be defining factors.
We also show how the FRS gets hampered with an increasing $d$ based on the initial entropy. Notice how the LLaMA models suffer the highest drop in FRS (0.174-0.174 for 3b, and 0.149-0.149 for 8b), meaning that they were uncertain, on average, about the initial factual knowledge, confirmed in Figure˜1 where LLaMA has the highest accuracy drop across the board. To explore this further, Figure˜4 illustrates FRS with $d=1$ across all models on the SQuAD dataset. Notably, Qwen-14b achieves higher FRS scores even when its initial entropy is lower than that of GPT-4o-mini, whereas the latter’s scores are more concentrated toward the lower entropy range. Additionally, GPT-4o-mini exhibits a more skewed probability density distribution, with $\text{FRS}\rightarrow 1$, suggesting that while it may be highly confident in certain cases, its robustness is not consistently superior across all factual scenarios (see Appendix˜G). These findings highlight the need for a deeper examination into how architectures, training methodologies, and knowledge retention influence factual robustness.
@@ 5.3 Most and Least Robust Facts
We further analyze the specific question types that are the most and least robust within the models. For this, we choose to fine-tune a lightweight classifier, DistilBERT (sanh2019distilbert), on the TREC question type dataset (li2002learning) to classify the categories Numerical, Location, Entity and Human, for which we reach 96,6% accuracy on the test set. [^4] We report details about the dataset and the training procedure in Appendix H. We then collect the top 1000 highest and lowest FRS score samples across all datasets, for each model. For both the top and bottom samples, we compute the proportion of each question type relative to the total number of samples of that type. Figure˜5 shows the ratio of the most robust question types in percentage for each model. For example, for LLaMA-3B, 64.8% of numerical type questions are among the most robust (while $100\%-64.8\%=35.2\%$ are part of the least robust). Answers associated with numerical and location related questions tend to be the most robust across models, reaching up to 72.9% and 64.9%, respectively. The least robust answers belong to the Human entity type, which means questions asking about the name of a person or a group. We show that all models recall names in a less robust way, meaning that those are most prone to be produced with high entropy and/or are shifting to a wrong answer at low temperature levels. We hypothesize that this is due to the nature of expected answers: for example, questions of the numerical type most often call for a single digit as answer, e.g., a year. Here, the LLM has less individual tokens to produce, and hence less chance of introducing an error, than for example with the name of a person, which is more lengthy. It is furthermore interesting to note the difference between model sizes: while the bigger models perform highest on numerical type questions, they underperform smaller models like LLaMA-3B on human type questions.